find the term independent of x in binomial expression


(x4+1/x)10


(3x2-1/3x)9


(x-1/x2)3n

Grade:

1 Answers

Aman Bansal
11 years ago

Dear Deepika,

We know that any algebraic expression with two variables is called a binomial.

 Observe the following powers of the binomial (x + y).

     (x + y)0 = 1 

     (x + y)1 = x + y 

     (x + y)2 = x2 + 2xy + y2 

     (x + y)3 = x3 + 3x2y + 3xy2 + y3 

     (x + y)4 = (x + y)3(x + y) 

                 = (x3 + 3x2y + 3xy2 + y3) (x + y) 

                 = x4 + 3x3y + 3x2y2 + xy3 + x3y + 3x2y2 + 3xy3+ y4

     (x + y)5 = x5 + 5x4y + 10x3y2 + 10x2y3 + 5xy4 + y5

 

     What do you notice in the above expansion of various powers of (x + y)?

 
  1. The number of terms in the expansion is one more than the exponent.

  2. In each expansion:
  1. The exponent of the first term is same as the exponent of the binomial. The exponent of y in the first term is zero. 

  2. Subsequently, in each successive term, the exponent of x decreases by 1 with a simultaneous increase of 1 in the exponent of y. 

  3. The sum of the exponents of x and y in each term is equal to the exponent of the binomial. 

  4. The exponent of x in the last term is zero and that of y is equal to the exponent of the binomial.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert